A 5.0 cm3 solution of H2O2 liberates 0.508 g of iodine from an acidified KI solution. Calculate the strength of H2O2 solution in terms of volume strength at STP.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. The redox reaction involved is : H2O2 + 2I- + 2H+ → 2H2O + I2 If M is molarity of H2O2 solution, then 5M = (0.508 × 1000)/254 (∵ 1 mole H2O2 ≡ 1 mole I2) ⇒ M = 0.4 Also, n-factor of H2O2 is 2, therefore normality of H2O2 solution is 0.8 N. ⇒ Volume strength = Normality × 5.6 = 0.8 × 5.6 = 4.48 V